I need to solve these two problems by completing the square but I'm not exactly sure how to do it.
x^2 - 3x - 15 = -5
x^2 - 17x = -10x - 6

Answer by vksarvepalli(154)

x^2 - 3x - 15 = -5
=> x^2 - 3x - 10 = 0
=> x^2 - 5x + 2x - 10 = 0
=> x(x-5)+2(x-5)=0
=> (x-5)(x+2)=0
so x = 5 , -2


x^2 - 17x = -10x - 6
x^2 - 7x + 6 = 0
=> x^2 - x - 6x + 6 = 0
=> x(x-1)-6(x-1)=0
=> (x-1)(x-6)=0
so x = 1, 6